Question 7 1 pts You could also estimate a regression with arrest likelihood as an outcome when the dummy variable for treatment as the regressor. (This regression estimates what is commonly called the intent to treat effect). What would the coefficient on treatment be for that regression? arresti-α0 + α!treatment
Answer #1

The coefficient on treatment in the regression would be \alpha 1 which is the slope differential .i.e. when treatment dummy is 1 the slope is \alpha 0+\alpha 1 compared to when the treatment dummy takes value zero; the slope is \alpha 0.